Keep 1 to 9 in order and group to equal 100
Answer by nyc_function(2741)

To group digits 1 to 9 means to place small groups digits in parentheses whose sum will give you 100.

SAMPLE:
Let's say I was given digits 1 - 5 and given the same directions to find a sum of 15.
This is what I would do:
(1+2+3)+(4+5) = 6 + 9 = 15
This is exactly what you must do in orde to find your sum of 100.
Understand? This is what you must do with your question.